# HiLo: Detailed and Robust 3D Clothed Human Reconstruction with High-and Low-Frequency Information of Parametric Models

### v1

@@ 3.1 Progressive Growing High-Frequency SDF
Conventional high-frequency SDF. To improve the ability to represent complicated 3D shapes robustly, a straightforward way is to apply periodic functions $\mathcal{H}$ such as sine and cosine [44] to extract high-frequency signals on SDF of each sampled point via
@@ B.2 Details on Spatial Interaction MLP
As shown in Fig. A, take the $1$-th layer of $\phi_{si}$ as an example, the attention module $\mathcal{A}^{1}$ takes in the $\mathcal{M}^{3D}_{\mathcal{V}}$ and $\mathbf{F}_{c}^{(1)}$ and output a spatial interaction feature map $\mathbf{F}_{I}^{(1)}$. Specifically, We first extract a global spatial features $\mathbf{F}_{g}^{(j)}$ of the $\mathcal{M}_{v}^{3D}$ via a 3D Convolution block and a sigmoid function. We achieve the spatial interaction process of different voxels through the equation $\mathbf{F}_{c}^{(j)}\times\mathbf{F}_{g}^{(j)}\rightarrow\mathbf{F}_{I}^{(j)}$. After obtaining the $\mathbf{F}_{I}^{(1)}$, we fed it to the first full-connected layer $T^{(1)}$ to obtain $\mathbf{F}^{2}_{c}$.
@@ C More Experimental Details
1 Obtaining parametric body model SMPL-X $\mathcal{M}$ with $E_{p}(\mathcal{I})$.
2 With $\mathcal{M}$, obtaining the global voxel grid $\mathcal{M}_{v}^{3D}$ using $f_{3D}(\mathcal{V}(\mathcal{M}))$.
@@
4       Generating SDF $s$ w.r.t. $\mathbf{p}_{i}$ using Eqn. 1.Using $\mathcal{H}(\cdot~{};\beta)$ to enhance the SDF $s$ resulting in point-wise progressive high-frequency SDF $\mathcal{H}(s;\beta)$.
@@ C.4.1 Revisit Existing Methods
$\mathrm{PIFu}\mathrel{\mathop{:}}\phi(f_{\text{2D}}(\mathcal{I})(\mathbf{p})),\mathbf{p}_{z})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (C)
@@
$\mathrm{PaMIR}\mathrel{\mathop{:}}\phi((f_{\text{2D}}(\mathcal{I})({\mathbf{p}})),\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (D)
@@
$\mathrm{ICON}\mathrel{\mathop{:}}\phi(s(\mathbf{p}),\mathcal{F}_{\mathrm{n}})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (E)
@@
$\mathrm{D\text{-}IF}\mathrel{\mathop{:}}~{}$ $\hat{{\mathcal{O}}}_{f}=\hat{{\mathcal{O}}}_{c}+\phi_{r}(\hat{{\mathcal{O}}}_{c}\oplus{\mathcal{F}}_{7\mathrm{D}}\oplus P_{\varphi}({\mathcal{F}}_{7\mathrm{D}}))$ (F)
@@ C.5 Variant Methods
$\mathrm{ICON}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}\mathrel{\mathop{:}}\phi(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (G)
$\mathrm{D\text{-}IF}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}\mathrel{\mathop{:}}\phi({\mathcal{F}}_{7\mathrm{D}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))+\phi_{r}(\hat{{\mathcal{O}}}_{c}\oplus{\mathcal{F}}_{7\mathrm{D}}\oplus P_{\varphi}({\mathcal{F}}_{7\mathrm{D}}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{M}_{v}^{3D}(p)}\mathrel{\mathop{:}}\phi_{si}(\mathcal{H}(s;\beta),\mathcal{F}_{\mathrm{n}})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}_{s}(p;\beta)}\mathrel{\mathop{:}}\phi_{si}(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w}~{}\mathcal{H}_{s}(p)}\mathrel{\mathop{:}}\phi_{si}({\mathcal{H}}(s),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{\phi}_{si}}\mathrel{\mathop{:}}\phi(\mathcal{H}(s;\beta),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}(s;\beta)~{}\mathrm{w/o}~{}\mathcal{\phi}_{si}}\mathrel{\mathop{:}}\phi(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
@@ E More Experiments
SMPL-X Model. Skinned Multi-Person Linear-Expressive model (SMPL-X) [38] represents human body shapes and poses in a compact and parametric manner. The core idea behind SMPL is to use a linear combination of body shape parameters and joint rotations to represent a 3D human body model with $N{=}10475$ vertices and $K{=}54$ joints. Specifically, SMPL-X is defined by $\mathcal{M}(\theta,\beta,\psi)\mathrel{\mathop{:}}\mathbb{R}^{|\theta|\times|\beta|\times|\psi|}\to\mathbb{R}^{3N}$, where $\theta\in\mathbb{R}^{3(K+1)}$ represents the pose parameter, $\beta\in\mathbb{R}^{|\theta|}$ is the shape parameter, and $\psi$ denotes the facial expression parameters, and $K$ denotes the number of body joints in addition to a joint for global rotation. By adjusting $\theta$, $\beta$, $\psi$, SMPL-X is able to represent a wide variety of human body shapes and poses. See [38] for more details.
@@ F.1 Transfer Sketch to 3D model
Since our HiLo is robust to in-the-wild images, we are able to put it to more applications. We show in Fig. B that our HiLo is able to transfer a sketch image of a clothed human into a 3D model with the help of ControlNet [56]. Specifically, we collect sketch images from Pinterest and use ControlNet to transfer the images to RGB images. The RGB images are then fed to our HiLo to reconstruct 3D model of the corresponding human.

### v2

@@ 3.1 Progressive Growing High-Frequency SDF
Conventional high-frequency SDF. To improve the ability to represent complicated 3D shapes robustly, a straightforward way is to apply periodic functions $\mathcal{H}$ such as sine and cosine [47, 66] to extract high-frequency signals on SDF of each sampled point via
@@ B.2 Details on Spatial Interaction MLP
As shown in Fig. A, take the $1$-th layer of $\phi_{si}$ as an example, we use attention module [20, 25] $\mathcal{A}^{1}$ that takes in the $\mathcal{M}^{3D}_{\mathcal{V}}$ and $\mathbf{F}_{c}^{(1)}$ and output a spatial interaction feature map $\mathbf{F}_{I}^{(1)}$. Specifically, We first extract a global spatial features $\mathbf{F}_{g}^{(j)}$ of the $\mathcal{M}_{v}^{3D}$ via a 3D Convolution block and a sigmoid function. We achieve the spatial interaction process of different voxels through the equation $\mathbf{F}_{c}^{(j)}\times\mathbf{F}_{g}^{(j)}\rightarrow\mathbf{F}_{I}^{(j)}$. After obtaining the $\mathbf{F}_{I}^{(1)}$, we fed it to the first full-connected layer $T^{(1)}$ to obtain $\mathbf{F}^{2}_{c}$.
@@ C More Experimental Details
1
2Obtaining parametric body model SMPL-X $\mathcal{M}$ with $E_{p}(\mathcal{I})$.
3
4With $\mathcal{M}$, obtaining the global voxel grid $\mathcal{M}_{v}^{3D}$ using $f_{3D}(\mathcal{V}(\mathcal{M}))$.
@@
6       Generating SDF $s$ w.r.t. $\mathbf{p}_{i}$ using Eqn. 1.
7      Using $\mathcal{H}(\cdot~{};\beta)$ to enhance the SDF $s$ resulting in point-wise progressive high-frequency SDF $\mathcal{H}(s;\beta)$.
8
@@ C.4.1 Revisit Existing Methods
$\mathrm{PIFu}:\phi(f_{\text{2D}}(\mathcal{I})(\mathbf{p})),\mathbf{p}_{z})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (C)
@@
$\mathrm{PaMIR}:\phi((f_{\text{2D}}(\mathcal{I})({\mathbf{p}})),\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (D)
@@
$\mathrm{ICON}:\phi(s(\mathbf{p}),\mathcal{F}_{\mathrm{n}})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (E)
@@
$\mathrm{D\text{-}IF}:{}$ $\hat{{\mathcal{O}}}_{f}=\hat{{\mathcal{O}}}_{c}+\phi_{r}(\hat{{\mathcal{O}}}_{c}\oplus{\mathcal{F}}_{7\mathrm{D}}\oplus P_{\varphi}({\mathcal{F}}_{7\mathrm{D}}))$ (F)
@@ C.5 Variant Methods
$\mathrm{ICON}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}:\phi(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (G)
$\mathrm{D\text{-}IF}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}:\phi({\mathcal{F}}_{7\mathrm{D}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))+\phi_{r}(\hat{{\mathcal{O}}}_{c}\oplus{\mathcal{F}}_{7\mathrm{D}}\oplus P_{\varphi}({\mathcal{F}}_{7\mathrm{D}}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{M}_{v}^{3D}(p)}:\phi_{si}(\mathcal{H}(s;\beta),\mathcal{F}_{\mathrm{n}})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}_{s}(p;\beta)}:\phi_{si}(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w}~{}\mathcal{H}_{s}(p)}:\phi_{si}({\mathcal{H}}(s),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{\phi}_{si}}:\phi(\mathcal{H}(s;\beta),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}(s;\beta)~{}\mathrm{w/o}~{}\mathcal{\phi}_{si}}:\phi(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
@@ E More Experiments
SMPL-X Model. Skinned Multi-Person Linear-Expressive model (SMPL-X) [41] represents human body shapes and poses in a compact and parametric manner. The core idea behind SMPL is to use a linear combination of body shape parameters and joint rotations to represent a 3D human body model with $N{=}10475$ vertices and $K{=}54$ joints. Specifically, SMPL-X is defined by $\mathcal{M}(\theta,\beta,\psi):\mathbb{R}^{|\theta|\times|\beta|\times|\psi|}\to\mathbb{R}^{3N}$, where $\theta\in\mathbb{R}^{3(K+1)}$ represents the pose parameter, $\beta\in\mathbb{R}^{|\theta|}$ is the shape parameter, and $\psi$ denotes the facial expression parameters, and $K$ denotes the number of body joints in addition to a joint for global rotation. By adjusting $\theta$, $\beta$, $\psi$, SMPL-X is able to represent a wide variety of human body shapes and poses. See [41] for more details.
@@ F.1 Transfer Sketch to 3D model
Since our HiLo is robust to in-the-wild images [56, 31], we are able to put it to more applications. We show in Fig. B that our HiLo is able to transfer a sketch image of a clothed human into a 3D model with the help of ControlNet [60]. Specifically, we collect sketch images from Pinterest and use ControlNet to transfer the images to RGB images. The RGB images are then fed to our HiLo to reconstruct 3D model of the corresponding human.